OpenStudy (anonymous):

3x-y=26 -2x-y=-24 solve please?

OpenStudy (mathmate):

Subtract the second equation from the first to get 5x=50 => x=10 Can you take it from here?

OpenStudy (anonymous):

no :/

OpenStudy (mathmate):

3x-y=26 now we know that x=10, so substitute 10 for x to get 3(10)-y=26 30-y=26 What would y be?

OpenStudy (anonymous):

4?